40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 12 of 62  Not logged in ELOG logo
ID Date Author Type Category Subject
  2576   Tue Oct 27 15:57:27 2020 shrutiDailyProgressLab Workpsoma locking

Today I modified the optical setup with the aim of obtaining the beat between the two diode lasers for phase-locking.

I added pick-off polarizing beamsplitters with HWPs in each path for now (to be able to adjust their power) and mixed them at a 50/50 non-polarizing beam-splitter to eventually reach a Newfocus 1811 low noise PD.

I will add pictures and more details later.

 

  2575   Fri Oct 23 19:06:05 2020 gautamUpdateEquipment LoanTwo resonant RFPDs ---> 40m

I was in the Cryo lab between 1215-1230 this afternoon. I removed two resonant RFPDs from what was Johannes' setup (encircled in Attachment #1). I also brought a SR554 preamplifier to the 40m.

I was wearing the usual PPE (gloves, face mask) while I was in the lab.

Attachment 1: IMG_8903.JPG
IMG_8903.JPG
  2574   Thu Oct 15 22:15:11 2020 Ian MacMillanComputingSimulationModeringer Simulation

This is the spectrum coming off of the sample. there should be a peak at 1038Hz... but there isn't. And what is even weirder is that the spectrum analyzer that is built into Simulink shows a peak where I expect but when I do it here it doesn't show up. 

Update: I think I have found why there is a discrepancy between the two versions of the power spectrum. The spectrum analyzer in the Simulink model requires non-continuous data so you have to use a block to make your data discrete. The sampling rate of that block affects where the peak of the mode is seen. So it seems that that the mode seen in the previous post was just caused by making the data non-continuous.

 

Attachment 1: ModeRingerSpec.pdf
ModeRingerSpec.pdf
  2573   Thu Oct 8 00:31:11 2020 Ian MacMillanComputingSimulationModeringer Simulation

I got the script to run the simulation and added color to the diagram just to make it pretty :0

The next step is quantifying the error of the loop. My plan for this was to just calculate the Q from figure 3 of the moderinger paper. Then I can see if that value is consistent. Aaron suggested settling time and phase margin of the loop. So that is the next step (once I figure out how to do that)

Also I added the code to the Qryo github

Attachment 1: ModeRingerLoopDig.pdf
ModeRingerLoopDig.pdf
Attachment 2: ModeRingerSim.zip
  2572   Tue Sep 29 11:00:53 2020 aaronComputingSimulationModeringer Simulation

Beautiful! Want to push this to the repo under git large file storage?

Quote:

The step function represents the excitation of the sample. Ideally, it would excite to the setpoint and stay there but for some reason, it is jumping way past the point before returning. By messing with the gain and the frequency of the low pass filter I could get a variety of results the best is shown below at 50 gain and 40 rad/s.

The overshoot is interesting! To understand the loop shaping, I suggest checking out Gardner's Phaselock Techniques or Astrom and Murray's Feedback SystemsThey both have sections on optimal PID controller design (at least Astrom and Murray do). You can make a pole-zero plot to help choose the location of poles and zeros in your loop shaping filter (the discrete zero-pole TF we added, after the gain).

 

  2571   Tue Sep 29 02:02:54 2020 Ian MacMillanComputingSimulationModeringer Simulation

Working with Aaron's suggestions (In the previous post) we got the mode ringer to converge. Previously the loop would continue to excite the sample to infinity but by fixing the following things we were able to get the step function to converge.

  • Added a discrete zero-pole transfer function after the 100 gain
  • Included the RMS block
  • Disabled zero-crossing detection in the saturation
  • Fixed the sign of the s^2 in the bottom of the plant TF (the i^2 included in s cancels the negative)

The step function represents the excitation of the sample. Ideally, it would excite to the setpoint and stay there but for some reason, it is jumping way past the point before returning. By messing with the gain and the frequency of the low pass filter I could get a variety of results the best is shown below at 50 gain and 40 rad/s.

The power spectrum shown is taken from the spectrum analyzer shown in the loop. It shows what I would expect with a peak around our mode frequency of 1038 Hz.

Next Steps:

  • Figure out how to run the model entirely from a script. (I got this partially working but it is not fully functional)
  • Fix the step function it still seems to be wild and we NEED to fix that before making meaningful progress on this project.
  • Add some sort of error estimation: quantify how good this loop is at measuring what we want.
    • How can we optimize our loop parameters to minimize this error?
  • Make it look pretty (black and white is gross)
Attachment 1: ModeRingerLoop.pdf
ModeRingerLoop.pdf
Attachment 2: StepPlot.pdf
StepPlot.pdf
Attachment 3: PowerSpectrum.pdf
PowerSpectrum.pdf
Attachment 4: ModeRingerSim.zip
  2570   Tue Sep 22 21:30:07 2020 aaronDailyProgressLab Worklab entry, supplies

notice of lab entry

Tue Sep 22 21:33:30 2020

inventory

I'm cleaning a bit, and gathering items not in use or in need of repair. They would make less mess in my office.

took an inventory of optics cleaning supplies, first aid kit, general cleaning supplies, wipes, etc. I found most were included in the first round inventory, but I took photos this time to convince my future self of object permanence. Will add to the wiki and update in a bit.

entered QIL, CTN

gloved up, shoe covers, went to QIL to check out the sprinklers and CTN to grab a GHz spectrum analyzer (HP8560E).

out: Wed Sep 23 00:20:03 2020

  2569   Tue Sep 15 12:29:41 2020 ranaLab InfrastructureHVACHEPA filters for PSOMA enclosure

I suggest just using what's used for th QIL table.

  2568   Thu Sep 10 10:21:34 2020 aaronLab InfrastructureHVACHEPA filters for PSOMA enclosure

Yesterday into today, I've been shopping for laminar flow HEPA fan filter modules for the PSOMA optical enclosure. I didn't find a lot of LIGO documentation listing specific filters, but here's what I've found online with some downselection on 'low vibration / high filter quality'. Please let me know If there's a company we often use, or if you can help direct this search at all. 

Another consideration is flow rate relative to our volume. I can do this calculation, but what is the particle density (eg at 0.3 um) we want to achieve inside the enclosure? I realize that I never got the particle counter recording despite it being on my list, so I'll try to do that remotely today. We need to know the ambient particle count and the clean volume in the enclosure.

Company

Part number / name

Price

Filter

Sound

Link for info

Notes

Terra Universal Smart Whisperflow Fan Filter Unit $1,111 HEPA (H10-100k) 40 dB @ 90 ft/min velocity @ face datasheet, product page The 'Smart' model has 9dB less noise than the standard, but costs about $500 more. Also comes in 'low profile' version.
Clean Air Products CAP118   HEPA -- product page requested quote
Thomas Scientific Fan filter module, "low noise" $1,760 HEPA   product page  
Price Industries LFDC series laminar flow diffusers       product page, datasheet  
Envirco Mac 10 original     51 dBA @ 90 FPM datasheet referenced in T2000382
A-J MFG SSLFHFD-FP "criti-clean" fan filter   HEPA 45dBA @ 90 FPM product page claims to be the lowest-height unit on the market (13.7''). The RFQ page looked fishy though, so I held off.
FungiFun (website)       link lists a number of other filter suppliers, by country

*all prices listed here are from publicly available pages

  2567   Wed Sep 2 16:46:39 2020 aaronThings to BuyDrawingsdimension-full optical layout

I updated the stage 1/2 optical layout to be more detailed after getting a sense of the sizes of things again last week. Even though this isn't how the table is currently set up, it might be good to accommodate future vacuum chambers in our earlier designs to minimize how much we need to move and realign optics. 

  • just for convenience, I gave all of the vacuum chambers the footprint of the gyroscope vacuum chambers. I chose this footprint rather than the cantilever cavity cryostat footprint to allow 45 deg aoi on the mach zehnder BSs (the cryostat windows are all parallel)
  • The chamber containing the output BS and BHD is cramped. If these all need to be in the same vacuum, we should get a larger vacuum; if the BHD and LO path can be moved to air that works too.
  • If output BS chamber is larger and input MMT allows, the entire MZ can be shifted to the right (north). 
  • Picked off LO from input BS AR surface
  • Need additional mode matching lenses into some of the photodiodes -- certainly into PL_RFPD, but I think also into the other PDs?
  • The beam to TRANS_DCPD is probably geometrically unphysical. 
  • The cavity length is limited by the size of the cryostat.
  • Perhaps want free-space EOM, EOAM, Faraday isolator?
  • Not sure I've dumped all beams properly, for example the reflection from BHD_PD1 and from BHD_BS AR surface.

I will update the PSOMA hardware inventory tomorrow to reflect the additional details in the new drawing. The updated diagram is available on git LFS, and the hardware inventory now reflects the diagram up through stage 1.

Attachment 1: psoma_stage2.pdf
psoma_stage2.pdf
  2566   Wed Sep 2 16:19:57 2020 shrutiUpdateLab Worklab entry, information gathering

Notice of lab entry: 20 Sep 2020   evening

Fiber modulators on the table :

1. Intensity modulators (BW: up to 12 GHz) MXAN-LN-10

2. EOM phase modulators (BW: up to 150 MHz) MPX-LN-01

Dimensions of vacuum cans mentioned in attachments.

Attachment 1: 7BA40146-4F86-4819-8D6C-FFDC8F246E50.jpeg
7BA40146-4F86-4819-8D6C-FFDC8F246E50.jpeg
Attachment 2: 129FE089-1E57-4E6D-9E77-F80F8E124CD1.jpeg
129FE089-1E57-4E6D-9E77-F80F8E124CD1.jpeg
  2565   Mon Aug 31 10:00:25 2020 aaronUpdateLab Workoptics setup

Here's the layout.

Some easy things that should be changed:

  • Transmission through the cavity end mirror should go to a steering mirror before PDA20CS
  • Should clamp the cables (eg power to PDA255) with something soft so they stay in place and aren't strained
  • Need to dump the reflection from transmon PD. Do we really need to dump transmission through mirrors with frosted backs?
Attachment 1: 200828_layout.jpg
200828_layout.jpg
  2564   Fri Aug 28 16:59:59 2020 aaronUpdateLab Workoptics setup

I measured the transmission of the Coastline 1m mirror at 180. ppm (S122C).

Alignment procedure while setting location of optics:

  1. use 2x irises to constrain a beam path at the locations of the eventual MZ input BS and the cavity BS.
  2. Place the first (R~1) flat mirror at 45 degrees centered on the beam line. Use this mirror to steer into an iris at the location of the cavity's curved mirror.
  3. Use steering optics to direct the eventual transmission beam into the trans mon PD (PDA 20CS)
  4. Place the curved mirror at near normal incidence. Use this mirror to steer the beam through irises at the location of the MZ output BS and cavity BS.
  5. Use steering mirrors to send the beam to the refl PD (PDA255).
  6. Place the cavity BS and steer the reflected beam to the refl PD.

Alignment procedure subsequently:

  1. Misalign the cavity BS
  2. Use the cavity mirrors to steer the circulated reflection into refl PD.
  3. Use cavity BS to align prompt reflection into refl PD.

 

  2563   Thu Aug 27 13:24:30 2020 aaronLab Infrastructurestuff happenslocked out of optomechanics cabinet

Not sure how, but none of the drawers of the blue optomechanics cabinet are opening. I don't have a key. Here's what happened

  1. I closed one drawer ("A") and opened the drawer below it ("B")
  2. Closed B
  3. Opened A and B simultaneously; I could not open only B, only A and B together. Close A and B.
  4. Open both A and B. Close only A.
  5. Close B
  6. None of the drawers now open, even A and B together. seems locked

Found someone who's had this problem before, might give it a try...

[update]

This worked, I used the metal meter stick to unlock the drawer.

[Stephen's updates]

--> note that link formatting breaks link for me, so here it is - https://www.practicalmachinist.com/vb/general/help-my-lista-locked-me-out-how-do-i-open-201606/

--> wrote up a similar experience with additional detail ENG_Labs/260

  2562   Wed Aug 26 12:12:22 2020 aaronUpdateLab Workoptics setup
  • aligned beam along the NS axis using two irises and the existing 2x f=100 lenses from the E path.
    • HWP->steering mirror-> 90-10 BS -> iris -> lens -> lens -> iris -> PD
    • all transmitted beams dumped
  • Mounted a 1'' x 1m FL Si mirror in a polaris mount. Made a ring cavity with that and 2x mirrors coated for 45 deg.

optics

Can anyone tell me the specs / history of some of the custom optics in cryo? I'm mounting the 1m Coastline mirror and will start with that in the PSOMA cavity.

  • Laseroptik 1'' mirrors. Batch # 19028Kel. Part(?) numbers L-13997, L-13998. (these are in red boxes cases w clear lid)
  • Coastline Optics 1.0''x0.25'' FS substrates. Coated for 1550nm at 45 degrees. (these are in stacks cylindrical clear cases, and include some witness and rejected samples)
  • Coastline optics 1'' silicon substrate, 1 meter focal length. SN 1.0-Si-1.0M
  • Photon Laseroptik 1'' FS substrates (~50), labeled 75S308926. Also labeled with a wedge angle or what looks like a focal length (no unit)
  • Two stacks from Gooch & Housego, taped in bubble wrap with the data sheet. These are coated for 1550nm, transmission ~1% S and ~5% P at 45 deg.
  2561   Tue Aug 25 13:42:40 2020 aaronUpdateLab Workoptics setup

setting up PSOMA beam path

I took some photos of the existing layout. I'll just take apart the E beam path, and leave the W path unchanged for now as reference.

I moved the E fiber output coupler closer to the edge of the table, to make this path easier to reach.

Hopped around on the laser hysterisis curve for a minute. To optimize the temperature,

  1. Started with TEC on near room temperature, then turned on the laser driver with the current set to maximize power on the meter (S122C)
  2. increased the temperature setpoint (decrease resistance setpoint) slowly until the laser power starts to decrease.
  3. Turn off the laser driver. Decrease temperature setpoint to just below the maximum power setpoint.
  4. Turn on the laser driver. Total power is now 4.6mW, compared to ~2mW with TEC off.
  2560   Tue Aug 25 13:23:49 2020 aaronThings to BuyGeneralcryo lab inventory

I started a cryo lab inventory that is separate from the PSOMA hardware inventory, and intended for stock items in the lab (optics, electronics, clamps, general safety and cleaning supplies, etc). It will be a work in progress. Both are accessible to anyone logged in to google drive with their ligo.org credentials.

Both are also linked on the PSOMA project wiki.

  2559   Mon Aug 24 13:22:32 2020 aaronUpdateLab WorkalaPSOMA

Did some mode matching, see the git.

  2558   Mon Aug 24 13:21:40 2020 aaronUpdateLab Worknotice of entry

Notice of lab entry

Date of entry: 24 Aug 2020

time of entry: 12pm
time of exit:  
 
Purpose: stage 1 cavity setup
  2557   Wed Aug 19 12:15:40 2020 aaronUpdateLab WorkNotice of entry

The cameras were unfortunately lost in the mail, but we can use my laptop or other camera. Ended up leaving to do a couple comsol things that needed completing today.

  2556   Wed Aug 19 12:11:47 2020 aaronUpdateLab WorkNotice of entry

Notice of lab entry

Date of entry: 18 Aug 2020

time of entry: 12pm
time of exit:  
 
Purpose: Identify some combination of optics that will let us mode match to our ring cavity, set up PDH locking with the existing ~33 MHz modulation.
  2555   Tue Aug 18 21:35:49 2020 aaronUpdateLab Worknotice of lab entry
I'm taking inventory of our optics supply, in between figuring out what's wrong with the temperature log channels.

The inventory is in Clickup, which is a new organizational tool I'm trying out. There' an easy csv export, so I can get it elsewhere if/when we want. We have a wide variety of lenses:

Focal length (mm) 25.4 38.1 100 103.22 154.82 200 225 250 258.04 309.65 412.86 500 850
Quantity @ 1550nm 1 4 2 7 2 2 6 2 1 3 12 1 2

I found fewer curved mirrors, but there were a couple.

Steps I took for the temperature sensor:

1. Tried to see what's the temperature by reading the current temperature in ndscope or dataviewer. In dataviewer, go to 'Signal' tab, and enter the channel name or find it on the list of slow channels. For ndscope...

nsdscope X1:AUX-LAB_TEMP_F 

No data appears on either. I restart cymac1, which seemed frozen, but still nothing.

2. What's going on with these channels?

$ caget X1:AUX-LAB_TEMP_F
X1:AUX-LAB_TEMP_F     -459.7
$ caget X1:AUX-LAB_TEMP_K
X1:AUX-LAB_TEMP_K     -0.0166677

Looks like the channel is reading zero.

3. I traced the cable from the particle counter and found that it sends data to cominaux, the common auxiliary machine for the lab.

ssh cominauxThat's a little surprising to me, because
cd ~/services/modbusIOC/
less CRYOXT.db

This is the database file that defines the channels on cominaux. I search for 'LAB_TEMP_F' and find the epics record for the temperature channels. The epics records are all "calc" records, and the temperature in Kelvin is taken from X1:AUX-ACROXT_AI_15. This corresponds to channel 15 of the acromag slow ADC.

$ caget X1:AUX-ACROXT_AI_15
X1:AUX-ACROXT_AI_15     -0.0003333

 That's starting to make sense, the cable from the particle counter didn't go to the acromag ADC. Starting from the ADC channel 15, I traced the cable back to what used to be the AD590 temperature transducer.

4. Where did the IC temperature sensor go? Searching the elog and my dusty memory... neither readily recalls where it went. Let's get another one, they are cheap and easy to use.

  2554   Tue Aug 18 15:05:33 2020 aaronUpdateLab Worknotice of lab entry

Notice of lab entry

Date of entry: 18 Aug 2020
time of entry: 2pm
time of exit:   11pm

Purpose: Inventory lens supply, identify some combination of optics that will let us mode match to our ring cavity. Picked up some books on silicon from the library.

  2553   Mon Aug 17 15:52:13 2020 aaronDailyProgressLab WorkNotice of lab entry

Measurements around table

  • Clearance of lights -- 100''
  • clearance of  beam -- 105''
  • distance between lights -- 6'
  • there's also a sprinkler head a few inches lower than the lights
  • distance from top of table to ground -- 35''
  • distance from bottom of table to ground -- 24''
  • distance between sprinklers -- 8'

Could probably move the table a few inches from the wall and make use of the space between the lights for the enclosure. There also isn't much room in the back corner in the NS direction, and we may want to shift in either direction. Orientation seems as good as it could be. The ceiling above PSOMA is lower than above cryo cavs.
Will mark up photos and post.

 

Attachment 1: Ceiling_lookingSouth.jpeg
Ceiling_lookingSouth.jpeg
Attachment 2: Ceiling_lookingNorth.jpeg
Ceiling_lookingNorth.jpeg
Attachment 3: Clearance.jpeg
Clearance.jpeg
  2552   Mon Aug 17 15:10:58 2020 aaronDailyProgressLab WorkNotice of lab entry

Notice of lab entry

Date of entry: 17 Aug 2020
time of entry: 3:30pm
time of exit:   
 
Purpose: Make sure unnecessary electronics in the lab are off, due to the CA energy Flex Alert. Take some photos of cryo cavities enclosure and filter to assist with new enclosure design.
  2551   Fri Aug 7 11:35:09 2020 aaronUpdateLab Worklab work

Entered the lab at 11:00 am. The lab is far too hot (78F) and humid (45%)! 

Inventoried available photodiodes in the cryo lab, on the PSOMA wiki.

Fri Aug 7 12:55:53 2020
Facilities came to check the fire alarm. I let them in with shoe covers, they held a magnet up to the fire alarm and left. Total operation took less than 5 min.
[lunch and a couple meetings 1 - 3]

Pretty tired honestly, I submitted an order for a few things we want in the lab, such as:

  • 2x "ultra quiet" fans to circulate air
  • longer ethernet cable for reaching far end of PSOMA table
  • headset web cam so we can take video while doing lab activities
  • more plastic storage bins
  • food and water for our emergency kit

The PSOMA optics table is from TMC vibration control (TMC 784-29701-01). I sent them an inquiry for the datasheet, and also asked whether they sell clean enclosures or have application notes that might help us put one together. Also talked with Arian Jadbabaie a bit about the Hutzler lab enclosures, and got some photos.

Exit ~430 pm

  2550   Wed Aug 5 19:03:56 2020 aaronUpdateLab WorkNotice of lab entry

Notice of lab entry

Date of entry: 7 Aug 2020
time of entry:  11:00 
time of exit:   
 
edit: postponed this work from 6 Aug to 7 Aug.

Purpose: Measure optical table dimensions, start to assess what we need for an enclosure. Set up EOM/EOAM to take some transfer functions. Set up realtime model. Figure out why particle counter isn't logging to frames. Turn on the HEPA for cryo cavs table.

  2549   Tue Aug 4 14:08:44 2020 shrutiDailyProgressLab WorkInitial experiments and prep: Beam profile

Notice of lab entry: I will be at the cryo lab for a few hours today (04 Aug 20) afternoon


Update:

Beam profile results [razor-blade method]

At a few locations along the beam I measured the vertical and horizontal beam radii by measuring the power at different positions of blade edge across the beam. The power measured at the photo-diode was fitted with  a + b \times {\rm erf}(\sqrt{2}(x-x_0)/w_z)) where w_z is the beam radius at the location along the beam and x was either the horizontal or vertical position of the blade.

Results:

The standard deviation as estimated from the fits are lower than \pm0.002 mm for all estimated radii.

But with the crude nature of the setup and not having the laser temperature stabilized (resulting in power drifts while taking readings), I guess that the error in each measurement is higher. The data in Attachment 3 also has recorded the resistance of the temperature sensor at different points while taking the data.

Beam radii
Distance [in.] Horizontal radius [mm] Vertical radius [mm]
5 \pm 0.1 0.120 0.121
6 \pm 0.1 0.146 0.150
7 \pm 0.1 0.175 0.173
8 \pm 0.1 0.203 0.200
9 \pm 0.1 0.231 0.229

 

- Attachment 1 has the updated setup with the clamped translation stage

- Attachment 2 shows the measured points with the error function fits. The offsets between the different curves along the x-axis are arbitrary.

- Attachment 3 shows a linear fit and an estimated divergence angle. I've assumed that all points I measured are outside the Rayleigh range, i.e., away the beam waist. The caption '5H' means that this set of data was taken 5 inches away from the output coupler (fibre-to-free space) moving the blade in the horizontal direction.

- Attachment 4 has all the data and jupyter notebook with analysis

- Attachment 5 shows the beam spot on the view card

Attachment 1: IMG_0268.jpg
IMG_0268.jpg
Attachment 2: Erf.pdf
Erf.pdf
Attachment 3: Div.pdf
Div.pdf
Attachment 4: beamprofdata.zip
Attachment 5: IMG_0258.jpg
IMG_0258.jpg
  2548   Mon Aug 3 13:10:46 2020 shrutiDailyProgressLab WorkInitial experiments and prep

Notice of lab entry: I will be at the cryo this (03 Aug 20) afternoon

Lights fixed, safety, cleanliness:

- On entering the lab I noticed that the work with the light fixtures was completed. Unfortunately I could not avert it or get it all covered in time; I had assumed that I would be contacted beforehand but was not. But, by inspecting the table I do not think it looked any dustier than before. For any such activity in the future after we’ve cleaned the optics, I will remember to get it covered beforehand.

- The particle counter did not seem to be saving any data so I’m unsure what the effect of this activity was. The 1 micron particle count:
When I entered (1 pm): 80
(6 pm) : 70
This change might probably be entirely attributed to the air scrubber.

- When Aaron and I chatted with Calum today about COVID safety, Calum pointed out that turning on any HEPA filters in the lab would reduce the time between single-person occupancy to almost 0 for our room, as it would serve as an additional air scrubber. I tried locating a switch on the cryo cavity table for the HEPA fans but could not locate it today.

Laser and beam profile related:

- After reading through bits of the laser operating manual and turning on the east cavity laser, I turned on the TEC (thermo-electric cooler) box [TED 200 C] with neither the TEC nor the servo on. I used this just to read off the resistance values of the sensor. With the current drive off it was 11.2 k-ohms [22 C].

- I waited until the sensor readings stabilized to changing at the level of a few ohms around 8.5 k-ohms [28 C] in a minute and began measuring the beam profile. (See Attachment 1 for setup)
I could not mount properly the only translation stage I found since the holes on it were smaller than 8-32, and moreover, I could not locate anything suitable that fits into the 1/4" holes on the table with this.

- Nonetheless, when I took some data and plotted it, it looked like an error function so I took more data. Will post plots later after fitting and analyzing. I may have to repeat this again.

- The sensor for the power measurement was S122C (Germanium) with a range of 700-1800 nm and 40 mW. After setting the power meter to 1550 nm, the measured power of the entire cross-section was ~2.3 mW.

- Also in the image in Attachment 1, the two optics (HWP and mirror) near the fiber-free space coupler, were previously in the path of the laser and now moved to either side of the beam. No other optics were moved from the previous set-up.

Attachment 1: IMG_633BBA57553C-1.jpeg
IMG_633BBA57553C-1.jpeg
  2547   Fri Jul 31 08:51:00 2020 shrutiDailyProgressLab Workbeam profile prep

Notice of lab entry: I will be at the cryo lab today (31 Jul 20) 9 am - 2 pm.

Purpose: Look around for components [power meter charger, power connectors for PDs, flashlights, etc], laser operating manual; find out specifications of available mirrors, RF PDs

Also, there seems to be only one laser temperature controller in the lab. The west cavity laser TEC port is not hooked up to anything. For the time being this is okay since we'd be using only one laser.

  2546   Thu Jul 30 09:13:14 2020 shrutiDailyProgressLab Workbeam profile prep

Notice of lab entry: I will be at the cryo lab today (30 Jul 20) 9am-5pm.

Purpose: Beam profile the Rio planex laser

- Turned on the laser that goes into the east cavity on. Turned on the Tenma supply, then the laser current driver  D1500207 labeled'E'. Saw a tiny, but bright, green spot on the detector card.

- Located razor blade and translation stage, for the beam profile measurement but not sure which power meter/ photodiode to use for 1550 nm. I plan to move the optics on the path of the east laser (to the east cavity) in order to do this measurement and later set up the initial ring cavity

- Aaron suggested I do another airflow measurement within the lab. Without the scrubber, there seems to be no measurable flow around either the Cryo-Q table or the cantilever table. I moved the scrubber into the lab near the workspace at the north side of the lab and let it run on medium speed. A few cm from the scrubber it the airflow rate is almost 5000 CFM, but it becomes <40 CFM even as close as the Cryo-Q table and again becomes negligible near the cantilever table.

- I set up the translation stage, moved relevant optics, but will continue to measurement later after I've researched and located a 1550 nm power meter

Two of the four overhead lights over the west (cantilever) table don't work anymore. Three other overhead lights are very dim/out.

 

 

  2545   Tue Jul 28 16:13:56 2020 aaronUpdateLab WorkAir flow measurement

Conclusions

  • I propose we continue scheduling cryo lab use with one individual using the lab per day, and announcing who will be in lab the day before.
    • If needed, this analysis suggests we could schedule two people to use the lab during the same day, but not at the same time. (of course, we are not yet cleared for shared use of the lab)
  • Use the air scrubber near the workstations, or wherever in the lab has stagnant air and prolonged use
  • Increase air circulation in the lab
    • I can order fans this week.
    • Can we open an additional port for air flow out (and possibly in) to the lab? What's behind the old cryo access in the NW corner, could we put a vent for out flow?
    • Change the HEPA filter
  • Before returning the anemometer, perhaps should measure air flow in the interior of the lab.
  • Check out the new cryo lab group and repo(s) on the github!

Cryo lab air flow

Liz dropped off an air scrubber (Medify Airx MA-40) and an anemometer (Digi-sense 20250-15) today. I'm using the instructions on the DCC to measure the air flow and assess the occupancy limits of the cryo lab. I calculate an acceptable amount of time between lab uses, and for two people to occupy the cryo lab simultaneously. For these calculations, I use a conservative threshold of P<1% for the acceptable probability that a second person becomes infected, given one infected lab occupant who sheds virus at 10 nL/min. I measured the dimensions of the lab at about 30x20x10' (l, w, h), for a 6000 ft^3 volume. I expect that's a high estimate, as it doesn't account for things like the awkward geometry of the staircase, volume of lab equipment, or stagnant air inside cabinets.

I measured the air flow at 20 points in a grid covering the surface of the HEPA filter. I also confirmed that there is no measurable air flow into the lab other than at the HEPA filter (eg, at the hole in the ceiling in the NW corner of the lab, or where the RF timing sígnal is sent through wall to Crackle lab). The average air flow through the HEPA filter is 200 CFM (cubic ft / min).
The flow is highest near the edges of the filter--does this indicate it is nearing the end of its life (it's been at least a year since replacement)? The flow is mostly above 200 CFM, but varies from 100-300 CFM across position and time.
 

Result

The first row is the minimum number of minutes to wait between when a (presumably infected) person exits the lab, and when a (presumably uninfected) person can enter the lab to maintain a <1% probability of new infection through aerosolized spread. The second row is the maximum number of person-hours the second and further (presumably uninfected) users can collectively spend in the lab with one presumably infected user, while still keeping a <1% probability of observing any areosolized spread. I don't have Spirou online currently, but I'll eventually put the calculation notebook in a shared folder accessible to the new cryo lab environment repo. Update: it's here.
 
Configuration cloth mask, no scrubber cloth mask, with scrubber N95+, no scrubber N95+,with Scrubber 
Minimum time between occupants (min) 17 7 0 0
Maximum person-hours of shared occupancy (person-hour) 0.004 0.008 0.08 0.15

These results are consistent with what I see for similar types of rooms in the LIGO spreadsheet.

Under an air flow model assuming perfect distribution of air from HVAC and HEPA scrubber throughout the lab, and very conservative requirements for probability of spreading COVID assuming a single lab user is infected, 

  1. Within 6 minutes of an infected user exiting lab, an uninfected person can enter and expect a tolerably low concentration of aerosolized virus.
  2. Within 10 minutes of one infected and one uninfected person working in the lab 6'+ apart with N95 or better masks, the viral load to an uninfected at-risk person may be intolerably high

Furthermore, the air in the lab Is not perfectly distributed. The air intake is several feet from the outflow, both near the door. The primary heat sources are the electronics racks along the N and NW sides of the room. The gradient from 71.1 F at the thermostat to 78 F at the cryocavs rack is uncomfortable to work in and bad for the electronics and optics. It's also too stagnant for the viral load conditions assumed in Evans P2000189 to apply. 

We are not currently cleared for 2+ people to use the lab. This analysis suggests to me that before doing so, we should improve the air flow conditions in the lab. And, even under optimal conditions we may minimize the total time with multiple people in the lab simultaneously.

Rio laser

I disconnected the oscillator, PDH boxes, and laser drivers from the power strip, and powered the strip with the Tenma supply. I had the ground and negative pins switched at first, careful of this -- ground is black, negative is green on this cable.

I plugged in the electronics, but only turned on DC power to the laser driver. I just left them on long enough to confirm a bright green (on viewcard) beam spot for both lasers -- we're in business!

Turned off the lasers, disinfected the common surfaces and objects.

PS, the anti-fog wipes work wonders. My goggles went from fogging in seconds to no fog at all.

  2544   Tue Jul 28 13:33:06 2020 aaronUpdateLab WorkRio laser

Notice of lab entry

[anticipated]

Date of entry: 28 July 2020
time of entry:  15:00 
time of exit:    19:30
 
Purpose: Turn on the Rio laser, find what we need for beam profiling, do some cleaning (mop, putting away stuff from the workbenches, etc)
  2543   Mon Jul 27 12:11:49 2020 aaronUpdateLab WorkSimplifying W optic bench

PSOMA hardware inventory

I located the materials for stage 1 PSOMA on the West optics table. I recorded what we have in the hardware inventory, and what we don't have is flagged for purchase. I start by cleaning up the electronics rack, removing anything I think is not in use.

Electronics rack

  • Both laser current drivers are D1500207
  • The E laser has a TED 200 C temp controller. The W laser is not plugged in to TEC, but there is a spare ITC 502 combination controller on the rack, and cryo cavs has a additional ITC 510.
  • There are three universal PDH servo boxes on the W rack: 1x D0901351 and 2x D1700219-v1. All of the boxes need lids and proper rack mount hardware.
    • Innards
      • Inside the *1351, the board matches the front panel DCC number (serial number 1437)
      • Both the *0219 have matching boards: D1700192-v2, D1700195-v2, D1799182-v2, D1700131-v2
    •  
  • Power supplies
    • The +- 15V power strip is currently unplugged. I don't see a low noise DC supply on this rack, but there is a Tenma 72-6615 on the ground. The closed-ring connector of the power strip is incompatible with this supply, so likely it was powered off of something else.
    • There's a NewFocus 1901 +- 15V supply on the W table itself. I think this should be located not on the table? The outputs are 0.1 A max to banana, plus 2x 0.3A max to a coax cable. The NewFocus was only supplying the PDs on the table. 
    • Cryo cavs has a dedicated rack for its power supplies (2x Sorensen DCS33-33E), separated from the laser control and PDH electronics. Is this an important choice, and should we also move our power supplies?
    • On the bottom of the rack, there's a Mech-Tronics NIM power supply. Seems out of use, only a PD power supply and readout board is plugged in, but this has no corresponding PD. 
    • I found on the elog that Chris moved the two Sorensens from the cryo cantilevers (W) rack to the cryoaux (central, rolling) rack after the supplies Johannes bought in 2017 died (fans not running). Do these usually die right after their 2 year warranty expires, or were we especially hard on them?
      • sidenote, great to see a potential solution to cryo lab timing woes, and that the overheated Sorensens were causing cominaux hard drive failures. Thanks Chris!
      • I've added 2x Sorensen supplies to our purchase list for PSOMA
    • For now, I moved all existing power supplies onto the electronics rack, and power the strip with the Tenma. In the future, we can discuss a remote location for power supplies, and will buy 2x additional Sorensens.
  • Misc
    • a bunch of old and open circuit boards, in cases compatible with the NIM supply. One power amp II; one PDH2 servo board (D1100996); 2 channel low noise amplifier (contains D1101396-v2 and another hand made board); a general filter board claiming to be from the gyro experiment. 
    • There were a few loose minicircuits components connected to the OCXO preamp (D1500064). These were providing 33.59 and 32.7 MHz modulation to Zach's experiments. I removed these from the preamp so we can give them a more permanent box later.

W optics bench

  • vacuum pump
    • I don't see the pumping station that used to be by the W optics bench. I don't remember using it elsewhere, and don't find any elog mentioning it being borrowed or moved. Has anyone seen this pump? It's one of the HiCube all-in-ones; it should have LIGO property tag C21832.
    • Similarly, what happened to the IR labs cryostat that housed the cantilever Q measurement (property tag)? I see maybe the glint of a cantilever in the short gryo vacuum can, and I have the cantilever Q HV driver and feedthrough, but no cryostat.
  • Photodiodes
    • PDA50B (Ge, 800-1800 nm)
    • 2x PDA255
    • Found 2 loose (unmounted) photodiodes wrapped in foil near the "optical contact razorblade test" setup (SE corner of W table).
    • NewFocus 1811, 1611
    • 1 home brewed QPD and breakout board, (PN on the circuit board is 10-00146 rev 1)
  • Put away or onto the workbench any loose optics, unused optomechanics, and most cables. Anything that wasn't a stock part I placed in a clear plastic box labeled 'Cryo Cantilevers' in the projects cabinet (2-part wooden cabinet on S wall). Stock parts went into the respective optics cabinets.

Misc

  • The lab temperature is still way too high. The temperature at the particle counter is reading 78, the thermostat still reads 71 F. The filter is not leaking substantially.j
  • Got the speaker working -- there was an aux adapter in the headphone port, so system was confused
  • Raymond requested a CF16 blank. I couldn't find one, but gave him a CF40-to-CF25 adapter that had a CF25 blank epoxied onto it, along with a CF16 to CF40 adapter and some gaskets. Here was the disinfection procedure for sending these items to QIL:
    1. Found a large plastic bag and disinfected inside and out with Phenomenal (TM) spray
    2. Disinfected the unpackaged UHV components with 70% isopropyl alcohol and a Kim wipe, wrapped them in foil, and placed in the disinfected bag
    3. Disinfect the outside of the bag with Phenomenal spray and leave outside the door to cryo for same-day pickup.
  • noted that pressure in cryo Q vacuum is 24 utorr after about a day valved off
  • Emailed Jordan and Liz about anemometer, to measure air flow in cryo. If we can compute the time for air to circulate through the cryo lab, we can reduce the length of time between lab entries (currently Shruti and I are leaving 1 day between our lab uses). 
  • Starting piling up seemingly unused components from W table onto the S workbench. These need a new home.
  • Found some historical documents, sticky notes from Zach. Took photos of these just in case (attached, later) and placed along with the miscellany on the workbench.
  • Checked with other W Bridge lab users before entering EE shop at 3:30pm. Disinfected surfaces I touched before and after use. I retrieved 3x 9V rechargeable batteries for the cryo lab multimeters.
  • Found a cardboard boxed fiber polarization controller (FPC032), which I placed in the NW optics cabinet sans box. Is this safe for storage (attachment 3)? What about things like Kim wipes, are those an acceptable amount of cardboard, or should they be eliminated?
  • Moved any cantilevers, etc into the silicon storage cabinet (NW corner cabinet)
  2542   Sun Jul 26 20:37:50 2020 aaronUpdate  

Notice of lab entry

[actual]

Date of entry: 27 July 2020
time of entry:  11:30 
time of exit:    19:00

 

Purpose: PSOMA hardware inventory, find out what we can turn on

  2541   Fri Jul 24 16:12:13 2020 aaronUpdateLab Workcryo open

entered W Bridge 264

I stopped by my office (W Bridge 364E) to pick up the new workstation computer and a mini monitor that might be handy. Stayed < 15 min and wore fresh gloves, mask, etc; also recorded entry by door.

I entered the lab around 3:15pm and wiped down the common surfaces. I sanitized and sorted my set of PPE from the box Liz dropped off -- thank you!

We should try not to bring cardboard Into the lab, and remove any cardboard we find. I got rid of a couple boxes.

lab temperature is high

The thermostat is doing its usual thing. It is set to 'cool,' but despite the temperature reading 71.2 F at the thermostat, the temperature at the particle counter is 80F and it feels warm. When I try to lower the set point, the thermostat requests a password. I found online that the default password for is 9995. The thermostat seems to accept the password but won't let me change the setpoint.

The rubber cement at one corner of the HEPA filter delaminated, but the edge held by the kim wipes set in place. I reapplied glue and fixed the tape more securely, and noticed improved air flow through the filter again. The lab temperature also improved a bit -- the leak seems directs cold air down to the floor, while the filter sends cool air across the top of the room.

misc reopening steps

These cover most of the recommended reopening steps. I'll do some general cleaning and organization, and make some new signage, then will call it 'reopen.' 

  • ran the sink hot and cold for about 10 min. There are no noticeable leaks or dampness under the sink, along the edges of the lab, or in any of the major cabinets in the lab.
  • Valved off the cryo Q vacuum and turned off the pump. Pressure reached 2e-3 torr in ~3 hours
  • strain relieved cables of spirou workstation, moved the station, ethernet, etc closer to PSOMA electronics rack
  • checked level of compressed He (1900 psi) and N2 (we don't have a cylinder). I will coordinate with Chub to get an N2 delivery
  • Added isopropyl, methanol, other solvents to the PSOMA inventory
  • Confirmed that our fire extinguisher still has pressure. Where is our eye wash and emergency shower station?  There's one on the basement level, is there one in the subbasement?
  • Not sure about sprinklers, gas, and vacuum utilities. I checked the utility with a yellow valve at NE corner of the lab, but opening the valve releases neither gas nor vacuum. I did not touch the sprinkler system, and did not test the compressed gas tubing (such as for the optics benches).
  • Laser 'in use' lkight at the lab entrance is functioning.
  • Checked that the hot plate (1) and variac cables were not under stress or damaged

exit lab Thu Jul 23 18:09:38 2020

  2540   Fri Jul 24 15:59:12 2020 aaronUpdateLab Workcryo open

Notice of lab entry

[anticipated]

Date of entry: 23 July 2020
time of entry:  15:15 PT
time of exit:    18:09 PT
 

Purpose: finish up the steps outlined above (2533)

[note: date is correct, I needed to remake the entries for July 23 but the content of 2541 is nearly identical. I had a backup.]

  2539   Tue Jul 21 14:03:33 2020 shrutiUpdateLab WorkAt cryo lab

Notice of lab entry: ~13:45. Expect to be here until around 17:00.

Liz dropped off PPE that was collected from the 40m (Thanks Liz!), this included two sets each of items as marked on the safety tab of PSOMA inventory doc. Also, see Attachment 1. I have removed one set of items and placed them on the tan cart that I use. Aaron's set is still in the box by the door; while I have worn gloves and tried not to touch those items, please assume that I have.

I think I have located the PD readout circuits requested by Gautam, which was in Johannes' setup. See Attachment 2,3,4; the DCC number of 2 checks out. I have placed them in a box in the area near the entrance marked by a red square. Other items labeled with '40m' are also in that box.

Attachment 1: 156CA9F4-5287-4222-9C45-5ACA006DFEDC.jpeg
156CA9F4-5287-4222-9C45-5ACA006DFEDC.jpeg
Attachment 2: 64F84876-1D78-427F-B272-5584D0D7AC2C.jpeg
64F84876-1D78-427F-B272-5584D0D7AC2C.jpeg
Attachment 3: B56E1C6B-C039-4E7E-85F1-AEFE8FA1BEAB.jpeg
B56E1C6B-C039-4E7E-85F1-AEFE8FA1BEAB.jpeg
Attachment 4: D540AD24-6D44-447C-B87F-B6CCD01D9521.jpeg
D540AD24-6D44-447C-B87F-B6CCD01D9521.jpeg
  2538   Mon Jul 20 08:59:02 2020 shrutiUpdateLab WorkAt cryo lab

Notice of lab entry:

After going over the checklist outlined in the previous elogs by Aaron, I completed everything relevant in the one labelled 'before', 'lab entry'.

Entry ~08:30, Will be here until ~17:30, with a break for lunch and a meeting (12:00-14:30)

Purpose: To give myself a lab tour, clean up the area around the workstation I'd be using, arrange and label tools, record part numbers of components for the PSOMA experiment

I also went in to the EE shop to get a solder station and related implements that I could find more than one of. I've placed this on the table to the south of the cryo Q experiment, after clearing up the space and placing the boxes below the table (see pictures).

Attachment 1: FFDA12DE-AD53-4751-96C5-7AA726F4BC51.jpeg
FFDA12DE-AD53-4751-96C5-7AA726F4BC51.jpeg
Attachment 2: 332F07DE-A250-44A6-BA28-5A0DF65207E9.jpeg
332F07DE-A250-44A6-BA28-5A0DF65207E9.jpeg
  2537   Fri Jul 17 13:15:03 2020 aaronUpdateLab WorkReopening cryo lab

TLDR: I sealed the leaks around the HEPA filter. I did an inventory of PPE and cleaning supplies in cryo, and separated two sets of personal PPE for Shruti and myself from the common supply. I did not receive PPE or cleaning supplies from Downs or 40m, but have inquired after both. Any anticipated work not explicitly mentioned here was not done; I tried to take care, but still strongly suggest that Shruti wipe down anything I placed in your personal supplies, just in case.

Next steps: See prior elog

 

After clearing the temperature check at home, still showing no symptoms, and starting the daily log in the reporting app, I head to West Bridge.

I entered the lab at 13:10 PT, using the N (olive walk) entrance to W Bridge. I am wearing a disposable mask, applied hand sanitizer on entering the lab, and don nitrile gloves. I locate the ISO and Kim wipes and wipe down

  • personal devices
  • glasses
  • keyboard, mouse, spirou workstation
  • door handles
  • Orange-brown lab chair
  • Common cabinet handles
  • label maker

Existing PPE and cleaning supplies in cryo

I located the following in the cryo lab, and am storing all PPE in the staging area at the entrance.

  • 4x laser glasses suitable for 1550nm. 2 are over-glasses, 2 are not, and I've labeled the glasses and storage cubby for myself and Shruti to have one of each.
  • 4x laser glasses for 1064. 2 are labeled '40m' and will be placed clean in the outbox for pickup with the circuit, the other 2 are also in labeled cubbies.
  • 1 pair of lN2 safety googles, 1 pair of ordinary sunglasses. These are not in regular use, and can be used occasionally with proper cleaning.
  • Limited supply of blue shoe covers (1 package)
  • 56 boxes of small dry Kim wipes, 4 box supply of large dry Kim wipes
  • 1 box of large nitrile gloves, nearly depleted
  • 2 Kimtech lens cleaning stations (34623; need to check whether this is just anti fog or also EPA approved sanitizing spray)
  • Various Tyvek coveralls (PN IC 182 BWH SM 00250C) size small to XL
  • 2 polyester lab coats (M, L)
  • Variety of Single-wrapped latex gloves, size 7, 8, 8.5, 9, 
  • hair nets
  • sleeve covers
  • 1 Willson face shield
  • disposable frocks size S 
  • limited supply of shop wipes
  • 1 package pre-wetted polyester wipes
  • dust catcher floor mats

Personal tools and PPE

Any tool I used that we have multiple copies of, I labeled "[aaron/shruti]" and set aside one copy for me and one for Shruti. My tools are on the blue cart holding Spirou. Shruti's are on the tan cart formerly holding screws and power supples (which I moved onto the remaining cart, the work bench, and the middle optics bench). Personal toolkit currently includes:

  • scissors
  • large flat head screwdriver
  • glasses cleaning station
  • dry kim wipes
  • small IPA spritz bottle 

Personal PPE are in clear bins labeled "Shruti" or "Aaron" and with the box contents. Currently each set includes

  • 2 disposable (but multi use) frocks
  • 2 disposable (but multi use) coveralls
  • 1 washable lab coat (where to launder?)
  • several pairs of AccuTech ultra-clean latex gloves
  • several hair nets
  • several blue shoe covers
  • several white sleeve covers

The remaining PPE are in clear boxes near the entrance labeled "common" and with the box contents. Extra AccuTech gloves and a few more coveralls are on the shelf below the HEPA filter.

Any tool with only one copy, I am noting here for a future order. I sanitized these tools before and after use. Common tools will remain in the tall blue Crafsman tool box.

  • label maker. The batteries were badly corroded (we usually run on power supply), I set them aside for disposal.
  • ladder

HEPA filter

The model of the HEPA filter is not visible from the exterior, and I will have to look up its cleaning schedule. The tape sealing the edges of the filter does not adhere to the textured paint surrounding part of the filter [attachment 1] . The smooth paint on the far wall appears to support adequate adhesion. To create a seal on the textured surface, I first cleared the supplies from the shelf under the filter and pealed back the non-adhering tape. I applied rubber cement to the sticky side of the tape, and held the tape in position for drying using the small boxes of Kim wipes from before. I applied rubber cement along the entire edge wherever there was textured paint, but not on the region with smooth paint. Air flow through the filter was noticeably increased following this procedure and seal was intact after a few hours, though I did not make a quantitative measurement. 

Stopped by Downs

at 1500 I exited the lab to use the restroom, eat, and pick up PPE from Downs 318. I swapped gloves on exit, and resanitized my hands after using the restroom in W Bridge basement, which is designated for our lab's use. After eating, I entered Downs and took the elevator to the 3rd floor, but was not able to access room 318 for PPE. I emailed Chub and Calum (listed contacts for Thomas lab) to coordinate a pickup at a later date. Specifically, I am requesting:

  • Nitrile gloves, M and S
  • 4 cloth face coverings, 2 for me and 2 for Shruti
  • Cleanroom masks (can also obtain from 40m PPE supply)
  • Beard covers (can also obtain from 40m PPE)
  • IPA wipes (some available in 40m supply)
  • disposable frocks size M (available at 40m)
  • 2x Disinfectant spray can
  • 2x hand sanitizer
  • 2x face shield
  • 2x no touch tools
  • 2x door adapter
  • 2x shop safety glasses
  • Signage
  • Borrow air flow meter if available 
  • anti-fog disinfecting wipes for glasses (is this the same as the Kimtech station?)

Misc

  • I found one function generator (SRS DS345) plugged into a power strip that was not connected to power. I've plugged it in to the wall and labelled it "check battery"

Exit

At 17:15, I begin wiping down the common tools and surfaces I used and exit at 17:25, upon which I disinfect hands again.

Attachment 1: tape_before.pdf
tape_before.pdf
Attachment 2: glasses.pdf
glasses.pdf
  2536   Thu Jul 16 12:09:18 2020 aaronNotesSafetyTemplate for logging lab entry / exit

Some of the above plans or guidance are now out of date. Here are some updates:

  • LIGO lab's research activities and restarting plan for PMA are outlined in a spreadsheet on the dcc.
  • There are also many additional or finalized training materials.
  • Most of the open questions raised above have been addressed
    • List of EPA approved cleaning products
    • Per the PPE and air flow requirements, since the Cryo lab will be a single person workspace we have no time constraint from COVID risk, and need only wear face covering when working solo. Other PPE requirements for clean optics still apply.
    • Shruti and I will be using the cryo lab on alternating days. Shruti will take Monday and Tuesday, Aaron will be Thursday and Friday. We will keep one unused day between our lab uses, at least for now.
    • We will set up a station with what we need from the electronics shop in the Cryo lab
    • LIGO has training and resources on best cleaning and disinfecting practice.
    • We can mostly use separate computer work stations, and keep personal tools at our own stations.
    • We will set up "in/out" boxes near the entry to cryo, so people can request supplies being stored there and pick them up after disinfecting. Information about shipping and receiving is on Sharepoint.
    • Rana has approved us to work in cryo 
  2535   Thu Jul 16 11:27:18 2020 aaronUpdateLab WorkReopening cryo lab

Notice of lab entry

[anticipated]
Date of entry: 17 July 2020
time of entry:  13:00 PT
time of exit:    18:00 PT
 
[actual]
Date of entry: 17 July 2020
time of entry:  13:10 PT
time of exit:    17:25 PT
 
Purpose: I will begin the reopening procedures outlined in previous elogs (2534, 2533). I'll also try to locate at least one Wilcoxon accelerometer and an M2 ISS readout board. I'll enter the EE shop to move one solder station from EE to cryo lab. 

That's probably all that's necessary for a typical lab entry, though more details always good. Here's what I'll be doing and a couple checklists for COVID procedures.

Planned activities
  1. Go to cryo lab and do a quick survey for PPE. I will pick up 2 sets of PPE, cleaning supplies, and signage from Downs 318 for cryo (one personal set per person). For list of PPE and other supplies, see the PSOMA hardware inventory (under Procurement on the wiki).
  2. Follow the applicable steps in the "back to lab" checklist in prior elog.
  3. ​ ​Put together personal tool cart and workstation, labeling
  4. Check / repair HEPA filter
    1. connection to wall
    2. check filter specification against requirements. air scrubbing reference. Calculator for acceptable person-time limits for shared rooms (currently not applicable, since cryo will be one person per day occupancy). Aerosol guidelines reference.
    3. See if particle counter still logging.
    4. Air flow meter measuring flow at filter, around filter, at the door, at the opening in the ceiling near NW corner.
  5. Set up in/out box bins for materials / tools entering and exiting cryo lab.
  6. Locate seismometer and circuit.
  7. Transport 1 solder station from EE shop to cryo
  8. Check inventory in cryo lab against the PSOMA hardware list
  9. Make sure cymac and the fe are running.
Before checklist

During checklist

  • Physically distanced at all times, following one-way signage around campus, wearing a face covering. I'll need to use elevators for a while, and will keep those single occupancy.
  • Wearing mask always and washing hands as appropriate.
  • On arrival, wipe down door knobs, keyboard, mouse, knobs, any frequently touched surface. 
  • No more than 1 person in the cryo lab per day, stay more then 6' from others.
  • Update other lab members of activity on the elog. The remote lab member is a 'virtual buddy'.

After checklist

  • complete elog, close work permit
  • Sanitize shared surfaces
  • Dof PPE except for face mask
  2534   Tue Jun 9 14:19:46 2020 aaronNotesSafetyTemplate for logging lab entry / exit

Here's my first attempt at a template / checklist for logging cryo lab entry / exit. In practice, perhaps the full checklist can be stored on a wiki page and also posted in the lab, rather than repeated many times on the elog. In that case, we can just include user name, time in and out, and purpose for lab use. Please provide changes, additions, comments, etc.

Cryo lab use log

User name: _____
Purpose: _____
Time in: _____
Time out: _____

Prerequisites to using the lab

By posting this elog, I confirm that I am eligible for returning to campus work by meeting the following:

  1. Have not shown flu- or cold-like symptoms, or been in close contact with anyone showing such symptoms, for at least 14 days.
  2. Watched Caltech's training video on returning to work safely.
  3. I am one of the allowed users in the LIGO instrument science group's plan approved by PMA.
  4. I have gotten the flu (and/or COVID) vaccine, unless medically prohibited from doing so
  5. I maintain a log of my activities on campus, including route travelled, buildings visited, and rooms visited, for each day that I am on campus.
  6. I wear a face covering at all times when on campus (is this right? is the requirement broader than on campus?)
  7. I know how to put on gloves, because I read the Glove Usage Guidelines attached (1)
  8. I have mastered the cleaning / sanitizing protocols outlined in attachment 2
  9. Usual lab safety measures apply, including long pants, closed-toe shoes, lab coats, and no food or beverage.

Lab entry checklist

By posting this elog, I confirm that on entering the cryo lab, I did all of the following, in order

  1. Check in with myself. I am not symptomatic and haven't come into contact with anyone showing flu- or cold-like symptoms. I meet all the above prerequisites for entering the lab.
  2. Check the elog for important updates -- is there any reason I cannot use the lab as planned?
  3. Apply hand sanitizer before entering the lab
  4. Donned personal protective equipment (PPE) in the following order
    1. Face covering / mask (worn before entering West Bridge)
    2. Gloves
    3. Hair covering
    4. Goggles (general or laser, depending on work)
    5. Lab coat
    6. Shoe covering
  5. Clean and sanitize shared work surfaces with 70% ethanol / water spray and wiping down.
    1. Doorknobs and light switches
    2. Keyboard, monitor, and mouse of shared computers
    3. Any chair I use
    4. Toolbox, cabinet, and drawer handles
    5. Work surfaces (benchtop on the south wall, computer desk, any other surface being worked on).
    6. Phone, personal electronic or other devices
    7. [list other shared surfaces necessary for today's work 
  6. Clean and sanitize my gloved hands, or change gloves

Lab exit checklist

By posting this elog, I confirm that on exiting the cryo lab I did all of the following, in order:

  1. Clean and sanitize my gloved hands, or change gloves
  2. Clean and sanitize shared work surfaces, including the ones listed above
  3. Remove personal protective equipment in reverse order from the above

Bathroom breaks checklist

By posting this elog, I confirm that on briefly leaving the cryo lab to use the restroom (or another room with users not from our lab), I did all of the following, in order

  1. Remove PPE (except for face mask) in the reverse order as above
  2. Put on a fresh pair of gloves before exiting the cryo lab
  3. Dispose of gloves before reentering the cryo lab
  4. Don PPE, in normal order as above

Cleaning and sanitizing protocol

The cleaning and sanitizing protocol is defined in attachment 2. Note that these apply to preventative measures to remove visible debris and reduce the number and growth of bacteria, virus, and fungi. Measures for disinfecting, or killing pathogens on a known contaminated surface, are defined elsewhere (see Caltech EHS).

  1. Clean the material to remove visible debris, dirt, or dust
    1. Use an EPA approved (dead link on the Caltech doc) cleaner, or either a 70% ethanol or 10% bleach solution.
    2. Use paper towels or wipes with cleaning solution
    3. Cleaning is performed whenever work surfaces or frequently touched surfaces are visibly dirty
  2. Sanitize the material to kill and reduce the growth of pathogens
    1. Wipe or spray surfaces with cleaning solution (70% ethanol, OR 10% bleach, OR another approved cleaning product), leaving the surface moist for at least one (1) minute
      1. For water sensitive equipment, use wipes
    2. Sanitation should be performed on any object used in the lab before and after use. Examples of these are:
      1. Bench tops
      2. Light switches, door handles
      3. Laboratory hand tools (wrenches, IR viewcards, etc)
      4. Waste container lids
      5. Equipment panels / controls, rack mounted electronics, cables
    3. Sanitation of commonly used surfaces (listed above) should take place before and after entry to the lab
  3. Discard used wipes, cleaning supplies, and gloves in regular waste.
Attachment 1: COVID-19_Glove_Usage_Guidance.docx.pdf
COVID-19_Glove_Usage_Guidance.docx.pdf COVID-19_Glove_Usage_Guidance.docx.pdf
Attachment 2: Work_Station_Sanitation_Protocol.docx.pdf
Work_Station_Sanitation_Protocol.docx.pdf Work_Station_Sanitation_Protocol.docx.pdf
  2533   Tue Jun 9 13:58:55 2020 aaronNotesSafetyPlan for reopened Cryo lab activities

Developing reopening plans

According to the plan for reconstituting lab activities in PMA, labs should plan with the following in mind. Open questions for cryo are in bold.

  • All work that can be done remotely should be done remotely.
  • Limit the total number of people with lab access, and the density of people at any time.
    • Establish designated individual work areas separated by ~10ft, depending on duration and type of activity. Use clear floor markings to delineate individual areas.
    • Establish traffic patterns (on-way aisles) to minimize contact.
    • How many people will be working In the lab? One is a natural start.
  • Establish shifts to minimize density, and break research groups into cohorts that have minimal overlap with each other.
    • Eg, 4 days in lab 10 days out; 3 days in 4 days out in two groups
    • Also, how will shared lab spaces like the electronics shop be used and scheduled?
  • Establish cleaning and disinfecting practices
    • High touch surfaces:
      • Doorknobs -- disinfect on entry and exit
      • tool boxes -- can we move to individual carts of tools?
      • Computers (keyboard, mouse) -- disinfect on entry and exit
      • Optics tables -- can we divide tables into individual zones, or give individuals their own table?
      • Chairs -- disinfect on entry and exit
      • Light switches -- disinfect on entry and exit
    • Disinfection strategy:
      • Cleaning with 70% iso (no higher or lower) from a spray (not squirt) bottle
  • Establish a plan for ordering and receiving supplies
    • One option could be to 'quarantine' deliveries in our personal, unused offices before bringing to lab. Cardboard boxes should never be brought to lab anyway.
  • Reassess lab safety hazards to address both COVID and effects of reduced numbers of people in labs
    • Maintain a log of all people in the lab (check-in and check-out). This can be done on the elog, I'll post a template elog entry shortly.
    • Require use of face coverings / masks, gloves, goggles, lab coat. We usually don't require lab coat in lab when not working on optics, but we should launder and use the lab coats on reopening. PPE should not be shared, and we already have plastic bins for individuals to store PPE between uses.
    • Require frequent handwashing and sanitation
    • Notify the safety office when a risky experiment is scheduled, and alert other lab members when working alone in the lab. (Eg, when venting an experiment, especially that contains HV).
    • Address safety hazard of unattended experiments. 
  • Establish a research continuity plan, including critical functions, communication plan, and data backup strategy. We have existing backups of things like our software and data which should be revisited.
  • Develop a communication plan. All lab activity should be documented on the elog, as always.
  • Identify a lab safety officer. This might be done at the LIGO-X level, rather than the cryo lab level?
  • Certify that the HVAC system complies with the standards attached (1). The cryo lab HEPA filter is currently leaky, needs repair.
  • Follow the back-to-lab checklist (attachment 2) for procedures on 'starting up' the lab again, including initial safety checks.
  • Additional guidance for developing plans for return to work is in attachment 3.

Back-to-Lab checklist

This section is an unpacking of attachment 2, with the specifics of the cryo lab in mind.

  1. Have a research group level 'back to work' plan developed. Have a schedule for starting up certain processes, equipment, etc, and avoid starting up alone (for safety). Take things slow.
  2. Perform a thorough walkthrough of the lab on first reentry, and note anything out of place, missing, damaged, leaking, etc. Record this In the elog, and make a plan for repair.
  3. Ensure the lab has adequate PPE for near-term use
  4. Ensure an adequate supply of hand soap, towels, 70% isopropyl, kim wipes, and other disinfecting equipment.
    1. We need to purchase a spray bottle, I don't know of one in cryo.
    2. We could designate the sole sink for hand washing, but may instead focus on hand sanitizer based washing and always wearing gloves in the lab. In any case, the sink should be flushed out until it doesn't run rusty on reopening.
  5. Complete an inventory of lab supplies
    1. I'll add that we should continue our campaign of reorganizing and labeling the cryo lab.
    2. Check levels of pressurized N2 (and liquid nitrogen, if needed).
    3. Discard expired chemicals, replenish supply of any depleted chemicals In addition to both high purity and low grade (70%) isopropyl.
    4. Replenish supply of PPE
    5. Supply of 1550nm optics, related optomechanics
    6. Inventory equipment used in old or unused experiments -- preserve experiments that are close to a functioning state, safely store experiments that are not
    7. Optics shelves
    8. Vacuum supply cabinet
    9. 'projects boxes' cabinet
    10. Toolbox
    11. Electronics racks
    12. Silicon sample and processing cabinet
  6. Verify all emergency equipment is accessible and functioning
    1. Eyewash station and safety shower in the hallways
    2. Sprinkler heads in lab
    3. Fire extinguisher
    4. Pull stations
  7. Check chemical and biological containers for damage, leaks, or pressure build up (we just have one flammables cabinet, and waste storage under the sink)
    1. Replace hazardous waste containers
    2. ensure compressed gas cylinders are chained and stored
    3. Maintain separation of non-compatible substances (oxidizers / flammables, acids / bases, etc)
    4. Review startup procedures for compressed gas cylinders
    5. Leak test compressed gas piping, especially going to floating optics benches
  8. Power up electrical equipment slowly and one at a time.
  9. Pour a small amount of water down dry traps and floor drains. Run all water faucets for a few minutes.
  10. Turn water back on slowly. Check water connections for leaks. Do not leave site right away, wait at least several minutes -- some connections may burst after few min. Immediately report all leaks.
  11. Check closed environments for accumulated condensation and/or mold.
  12. Check that utilities (house vacuum, H2, etc) are functioning.
  13. Equipment startup
    1. Review equipment manuals
    2. Verify "Laser in Use" lights, door interlocks, and other safety controls are functioning
    3. Verify heat sources do not have damaged cords. These include:
      1. 2x hot plates
      2. 1x variac heater, associated resistive straps
    4. Check vacuum equipment.
Attachment 1: Caltech_HVAC_System_Standards_Infomation.pdf
Caltech_HVAC_System_Standards_Infomation.pdf
Attachment 2: Researcher_Back-to-Lab_Checklist.docx.pdf
Researcher_Back-to-Lab_Checklist.docx.pdf Researcher_Back-to-Lab_Checklist.docx.pdf Researcher_Back-to-Lab_Checklist.docx.pdf
Attachment 3: PI_Return_to_Work_Lab_Planning_Guidance.docx.pdf
PI_Return_to_Work_Lab_Planning_Guidance.docx.pdf PI_Return_to_Work_Lab_Planning_Guidance.docx.pdf
  2532   Mon Jun 8 15:26:10 2020 aaronNotesSafetyPlan for reopened Cryo lab activities

Cryo reopening

Following discussion last week with Rana and Shruti, we would like to set up the early version on PSOMA in the Cryo lab. Working in cryo offers the following benefits:

  • Existing eyewear and controls for 1550nm laser. We identified 1550nm as a good wavelength choice, since it will eventually ease our transition from room temperature to cryogenic operation.
  • Access to cryo cantilevers apparatus
  • Possibility that cryo Q will move to the large cryostat down the hall in QIL (non-inverted cold plate)

What we need

I'm reviewing the existing COVID-19 guidelines from PMA and LIGO, and would like to develop a set of procedures for the people using the cryo lab to follow as we return to West Bridge. This includes:

  • Checklist of procedures for entering and exiting the lab (additional PPE, surfaces to clean, what to note on the elog)
  • Procedure for controlling lab access
    • for example, how many users may enter the lab per day or week, and how they should coordinate
    • One-way access to lab (hallways will become one-way corridors)
    • Daily record of contacts, symptoms, and work locations
  • What work needs to be done to prepare the lab for use? such as
    • Fixing the HEPA filter for HVAC
    • Restocking PPE
    • Post signage about new procedures
    • Sanitizing station near door
  • Other considerations?

Resources

Here are some of the resources available for answering these questions:

 

 

 

  2531   Tue May 5 13:05:36 2020 Ian MacMillanComputingSimulationModeringer Simulation

After an issue with Matlab on my computer, I am back to using python. This is the layout for the harmonic oscillator plant.

Attachment 1: ModeringerNotebook.ipynb
{
 "cells": [
  {
   "cell_type": "markdown",
   "metadata": {},
   "source": [
    "# Moderinger Notebook\n",
    "Ian MacMillan\n",
    "The goal of this notebook is to model the moderinger software."
   ]
... 146 more lines ...
  2530   Tue May 5 13:04:23 2020 aaronComputingSimulationModeringer Simulation

Looks pretty good. I suspect you're getting the diverging solution because the Q term in the denominator of your resonator TF is not imaginary. This means that the resonator has no damping, and any drive would indeed lead to diverging amplitude.

You might need to define the TF from the command line function, rather than in the simulink block; I think the block throws an error when you give it complex coefficients. This is surprising to me, so maybe I'm missing something.

Edit: That's not it, I just wasn't thinking in s-domain.

Here are a few other things I notice about this model that might be important:

  • The first BP filter is specified in Hz, rather than rad/s. This means you're inadvertently BPing well away from the plant's resonance.
  • You're computing the 'amplitude' by low passing the bandpassed signal. But what's done in the model is an RMS followed by lowpass. That is, the squared signal is scaled and lowpassed. Otherwise, you should get zero (BP near 1kHz followed by LP at 10s of Hz, with no nonlinear steps => 0)
  • You don't have the loop shaping filter, typically included in the PWRCTRL filter module. This contains the integrator needed to reach the correct setpoint (on long time scales, else you'll reach some offset from the setpoint) and an extra zero to increase the UGF.
  • The butterworth LP on the lower path has a pretty high cutoff, usually I set this at a few to 10 Hz.
  • The saturation gives a clean square wave iff there is a large gain just before, otherwise it gives a square wave with some rise/fall time where the original sine was crossing zero.
    • This is the source of at least some problems. Whenever I was increasing this gain, I found the output of differentiator decreased proportionally. I was also eventually getting some errors from the saturation block about detecting too many zero crossings. Turning off zero crossing detection at the saturation block solved both of these issues. In retrospect, this makes sense -- zero crossing detection is meant to smooth out discontinuities by decreasing the step size until they go away; this block is intentionally introducing such a discontinuity.
  • The laplace transform of an exponentially decaying sine has all positive real coefficients in the denominator. While the frequency-domain transfer function is what Rana posted above and has a (omega_0^2-omega^2) term, in the laplace domain it's different. I'd guess this is why you're geting a divergence.
  2529   Sun May 3 20:51:14 2020 Ian MacMillanComputingSimulationModeringer Simulation

Here is the model and the script for the moderinger as mentioned in the last post. the script should run it and plot the results. I haven't been able to test the script because of an unsolved MatLab issue on my laptop. 

Note: this model still has the problem of exciting the sample to infinity.

Attachment 1: Moderinger_V2.zip
  2528   Fri May 1 03:22:05 2020 Ian MacMillanHowToSimulationModeringer Simulation

I simplified the model and removed a few blocks to simplify the model mainly the RMS block and the block that just reduced the gain (it was pointless and mostly for messing with anyway). 

The idea with this model is to put in a set point then watch the model excite to that setpoint then hold the excitement of the disk to that setpoint. Then if I change that setpoint then the loop responds. 

I added a step that adds just a short signal to the loop for an instant to get things going. If the loop is working correctly this shouldn't change the outcome but should speed up the excitement because the sample always starts from a non-excited state. I have some readout boxes that see the excitement value of the sample and reports it.

What is currently happening when I turn on the simulation is that the excitement value either does one of two things: 1) It quickly goes to infinity. 2) It quickly goes to negative infinity. It takes less than a tenth of a second to do this, although my simulation is slowed down. Both negative infinity and infinity are basically the same thing: a very excited sample, where the negative infinity is a phase shift. From what I have found through experimentation:

1) The setting on the differentiator will change whether the sample goes to negative infinity or infinity based on whether the Maximum passband ripple (dB) is above or below 0.761

2) The speed of how fast it excites is based on the Butterworth filter Passband edge frequency (rad/s)

I'm not sure why these are true. I'm going to keep breaking it down and trying to find how it can correct itself and hold steady at the set point.

Since the simulation is perfectly repeatable (so long as you don't change any settings while you are running it) I haven't built it with code yet. It will become increasingly necessary when the model is able to hold an excited sample at a set point. Then we can pull raw data from it and increase the complexity of the plant.

Attachment 1: ModeringerSimV2Layout.pdf
ModeringerSimV2Layout.pdf
Attachment 2: simplemoderinger.slx
  2527   Mon Apr 27 19:15:32 2020 ranaHowTo Basic Simulink Tutorial

Mostly I've done it using linmod or linearize which are good for linear models. I haven't done it for nonlinear models in a long time, but this method might work.

ELOG V3.1.3-